# Counterexamples to two conjectures on Venn diagrams

### v3

@@ Abstract.
In 1984, Winkler conjectured that every simple Venn diagram with $n$ curves can be extended to a simple Venn diagram with $n+1$ curves. His conjecture is equivalent to the statement that the dual graph of any simple Venn diagram has a Hamilton cycle. In this work, we construct counterexamples to Winkler’s conjecture for all $n\geq 6$. As part of this proof, we computed all 3.430.404 simple Venn diagrams with $n=6$ curves (even their number was not previously known), among which we found 72 counterexamples.
While working on Winkler’s conjecture, Pruesse and Ruskey considered the (multi)graph whose vertices are the crossings of the curves of a Venn diagram and curve segments between consecutive crossings form the edges. They proved that this graph has a Hamilton cycle for every simple Venn diagram with $n$ curves, and conjectured that this also holds for non-simple diagrams. We construct counterexamples to this conjecture for all $n\geq 4$.
@@ 1. Introduction
A Venn diagram is simple if at most two of the $n$ curves intersect in any point: The three diagrams in the top row of Figure 1 are simple, whereas the three diagrams in the bottom row are non-simple. An easy application of Euler’s formula shows that a simple Venn diagram has exactly $22^{n}-$ 2 crossings, whereas non-simple diagrams may have much fewer. For humans, simple Venn diagrams tend to be easier to read than non-simple diagrams.
@@
A $k$-region in an $n$-Venn diagram is a region that lies inside of exactly $k$ of the $n$ curves (and outside of the remaining $n-k$ curves). A monotone Venn diagram is one in which for every $0<k<n$, every $k$-region is adjacent to both a $(k-1)$-region and a $(k+1)$-region. Bultena, Grünbaum and Ruskey [BGR99] proved that monotone diagrams are precisely the ones that can be drawn such that every curve has a convex shape. An exposed diagram is one in which the outer region is bounded by all $n$ curves. It is not hard to see that monotone implies exposed. Chilakamarri, Hamburger and Pippert [CHP96b] constructed $n$-Venn diagrams for every $n$ in which all regions are (combinatorial) triangles, quadrangles or pentagons, in particular, they are not exposed for $n\geq 6$.
@@ 1.1. Winkler’s conjecture
He reiterated the problem in his ‘Puzzled’ column [Win12] in the Communications of the ACM. The conjecture is also listed as the first open problem in Ruskey and Weston’s wonderful survey [RW97] on Venn diagrams.
Concerning small cases, the conjecture is easily verifiable for all simple Venn diagrams on $n\in\{3,4,5\}$ curves, simply because there are so few of them. Another class of Venn diagrams for which the conjecture is easily seen to be true are reducible diagrams. An $n$-Venn diagram is reducible if one of its curves can be removed so that the remaining $n-1$ curves form an $(n-1)$-Venn diagram. Otherwise the diagram is called irreducible. The diagram in Figure 1 (c) is irreducible. If a diagram is reducible, then the curve $C$ that can be removed splits every region of the resulting $(n-1)$-Venn diagram into two, and therefore this curve touches every region in the $n$-Venn diagram. We can thus add an $(n+1)$st curve by following the curve $C$ and subdividing every region into two, changing sides along $C$ exactly once along every segment between two consecutive crossings. Thus, the interesting instances of Conjecture 1 are irreducible diagrams.
@@ 1.1.1. Translation to a Hamilton cycle problem
Figure 2. (a) A 4-Venn diagram and its dual graph, a 4-Venn quadrangulation; (b) a Hamilton cycle in the quadrangulation; (c) the 5-Venn diagram obtained from (a) by adding the curve corresponding to the Hamilton cycle in (b).
@@
    It is a spanning subgraph of $Q_{n}$, i.e., all $2^{n}$ vertices are present. Specifically, the vertex in $Q(D)$ corresponding to a region of the diagram $D$ is the characteristic vector of the inside/outside relation with respect to each curve, where 0 and 1 represent outside and inside, respectively. [^1] When considering the diagram in the plane, then the inside and outside of a curve $C$ are the bounded and unbounded region of $\mathbb{R}^{2}\setminus C$, respectively.
@@
Extending the diagram $D$ by a single curve means that the new curve has to enter and leave every region of $D$ exactly once, splitting it into two, one of them inside and the other one outside the newly added curve. This corresponds to traversing a Hamilton cycle in the dual graph $Q(D)$, i.e., a cycle in the graph that visits every vertex exactly once; see Figure 2 (b)+(c).
@@
Figure 3. (Left) A counterexample to Winkler’s conjecture for $n=6$, namely a simple 6-Venn diagram that cannot be extended to a simple 7-Venn diagram by adding a suitable curve. Highlighted are 12 red regions that are surrounded by only 11 blue regions, which is a local obstruction for extendability. [^3] This drawing was produced using a SageMath program [Sch24] developed by Manfred Scheucher for his and Felsner’s work [FS21] on pseudocircle arrangements. We used this program with Scheucher’s permission, using as input the 6-Venn quadrangulation on the right. (Right) The dual graph of the diagram on the left, namely a 6-Venn quadrangulation $P$ that has no perfect matching and hence no Hamilton cycle. Edges are colored according to the colors of the curves. The highlighted subgraph $H$ contains a subset of 12 red vertices in one partition class and their 11 blue neighbors, witnessing the non-existence of a perfect matching. The green subgraph $L$ is the ladder used for extending the counterexample to all values of $n>6$, and all those extensions contain a copy of $H$.
1.1.2. The counterexamples
@@
As mentioned before, for $n\leq 5$ all simple $n$-Venn diagrams are extendable, so Theorem 2 is best possible in the sense that the counterexamples for $n=6$ are minimal. In fact, as part of our proof of Theorem 2 we computed all distinct 6-Venn diagrams and analyzed their properties, summarized in the next theorem, Figure 4 and Table 1. Specifically, we consider two diagrams the same if they differ only by mirroring and/or stereographic projection. Equivalently, we count their dual $n$-Venn quadrangulations up to graph isomorphism. Prior to this work, the number of simple $n$-Venn diagrams was only known up to $n\leq 5$. Specifically, for $n\leq 4$ there is only one, and for $n=5$ there are 20 [HP97, CHP00].
@@
The programs we used for computing all simple 6-Venn diagrams and the result files with all 6-Venn quadrangulations are available for download [Müt].
@@ 2. Proof of Theorem 2
Consider an $n$-Venn quadrangulation $P$ that contains a ladder $L$ and a subgraph $H$ with $L\cap H=\emptyset$. Then there is an $(n+1)$-Venn quadrangulation $P^{\prime}$ that contains a ladder $L^{\prime}$ and a copy $H^{\prime}$ of $H$ such that $L^{\prime}\cap H^{\prime}=\emptyset$.
@@
Let $x_{1},\ldots,x_{n}$ and $y_{1},\ldots,y_{n}$ be the vertices of the ladder $L$ in $P$ satisfying the definitions from before, and let $k\in[n]$ be the type of the edges $(x_{i},y_{i})$ for $i=1,\ldots,n$. We write $C$ for the outer cycle of $L$, i.e., for the graph $C:=L\setminus\{(x_{i},y_{i})\mid i=2,\ldots,n-1\}$. Note that in every 4-cycle of a Venn quadrangulation, there are edges of exactly two different types, and opposite edges have the same type. Consequently, both edges $(x_{i},x_{i+1})$ and $(y_{i},y_{i+1})$ of $L$ (and $C$) have the same type for $i=1,\ldots,n-1$, and because of the condition $x_{1}=0^{n}$ and $y_{n}=1^{n}$, every type $j\in[n]\setminus\{k\}$ appears for exactly one pair of opposite edges $(x_{i},x_{i+1})$ and $(y_{i},y_{i+1})$ for some $i=1,\ldots,n-1$. We write $C_{j}^{0}$ and $C_{j}^{1}$ for the two connected subgraphs of $C$ obtained from $C$ by removing both edges $(x_{i},x_{i+1})$ and $(y_{i},y_{i+1})$, such that the first one contains the vertex $x_{1}=0^{n}$ and the second one contains the vertex $y_{n}=1^{n}$.
We assume w.l.o.g. that the 4-cycle $(x_{1},x_{2},y_{2},y_{1})$ bounds the outer face of $P$. Let $P^{-}$ be the graph obtained from $P$ by removing the edges $(x_{i},y_{i})$ of the ladder $L$ for all $i=2,\ldots,n-1$. Let $\operatorname{rev}(P^{-})$ denote the plane graph obtained from $P^{-}$ by mirroring. Appending a 0-bit to all vertices of $P^{-}$ and a 1-bit to all vertices of $\operatorname{rev}(P^{-})$ gives the graphs $P^{-}0$ and $\operatorname{rev}(P^{-})1$. We let $P^{\prime}$ be the plane graph obtained by connecting $P^{-}0$ and $\operatorname{rev}(P^{-})1$ via $2n$ additional type-$(n+1)$ edges $(x_{j}0,x_{j}1)$ and $(y_{j}0,y_{j}1)$ for all $j=1,\ldots,n$. It is straightforward to check that $P^{\prime}$ satisfies conditions 13. In particular, the connectivity condition  3 is ensured via the subgraphs $C_{j}^{0}$ and $C_{j}^{1}$ for all bit positions $j\in[n]\setminus\{k\}$, and for the bit position $k$ it is ensured via the subgraphs $C0$ and $C1$.
@@ 3. Proof of Theorem 3
In this section we describe our C++ and SageMath programs that compute all 6-Venn quadrangulations. For the reader’s convenience, the code and data produced by them are available for download [Müt]. Specifically, the 6-Venn quadrangulations are provided in graph6 format, and in a compact binary adjacency list representation. The following explanations are valid for general $n$, but have a reasonable computation time only for $n\leq 6$. Table 1 provides the counts of Venn diagrams with various properties for all $n=1,\ldots,6$.
@@
Maybe surprisingly, the computational proof of Theorem 3 was the technically most demanding part of this work, even though in the end there are relatively few 6-Venn quadrangulations in total ($<10^{7}$ many). In fact, we had first found a counterexample to Winkler’s conjecture for $n=7$ (using different methods that involved SAT solvers) and a proof of Theorem 2 for $n\geq 7$, before finding the few counterexamples for $n=6$.
@@ 3.1. Dynamic programming on two halves of a Venn diagram
    It is a spanning subgraph of $Q_{n-1}$, i.e., all $2^{n-1}$ vertices are present.
@@ 3.1.1. Step (i): Computing all cycles $C$ of $Q_{n-1}$
In the following, we describe another observation that allows us to restrict the lengths of the boundary cycles to consider. As mentioned before, the total number of crossings in a simple $n$-Venn diagram $D$ is $22^{n}-$ 2. Clearly, this number equals the number of faces of $Q(D)$. The number of vertices of $Q(D)$ is $2^{n}$, and therefore Euler’s formula yields $42^{n+1}-$ 4 for the number of edges of $Q(D)$. These edges split into $n$ different matchings $M_{1},M_{2},\ldots,M_{n}$ according to their types, and we can therefore always guarantee that one matching $M_{i}$ has size $|M_{i}|\geq\frac{2^{n+1}-4}{n}$, and the fraction on the right should be rounded up to the next even integer, because the corresponding boundary cycle in the hypercube must have even integral length. For $n=6$ this yields the lower bound $|M_{i}|\geq 22$ for one of the matchings. By symmetry, we can thus restrict our search to cycles $C$ in $Q_{5}$ of lengths $22,24,26,28,30,32$. Note that if $C$ has the maximum length $2^{n-1}$, then it is a Hamilton cycle in $Q_{n-1}$, and the resulting $n$-Venn diagram $H$ obtained by pairing two compatible $C$-quadrangulations is reducible and thus extendable.

### v4

# Disproving two conjectures on the Hamiltonicity of Venn diagrams

@@ Abstract.
In 1984, Winkler conjectured that every simple Venn diagram with $n$ curves can be extended to a simple Venn diagram with $n+1$ curves. His conjecture is equivalent to the statement that the dual graph of any simple Venn diagram has a Hamilton cycle. In this work, we construct counterexamples to Winkler’s conjecture for all $n\geq 6$. As part of this proof, we computed all 3.430.404 simple Venn diagrams with $n=6$ curves (even their number was not previously known), among which we found 72 counterexamples. We also construct monotone Venn diagrams, i.e., diagrams that can be drawn with $n$ convex curves, and are not extendable, for all $n\geq 7$.
Furthermore, we also disprove another conjecture about the Hamiltonicity of the (primal) graph of a Venn diagram. Specifically, while working on Winkler’s conjecture, Pruesse and Ruskey proved that this graph has a Hamilton cycle for every simple Venn diagram with $n$ curves, and conjectured that this also holds for non-simple diagrams. We construct counterexamples to this conjecture for all $n\geq 4$.
@@ 1. Introduction
A Venn diagram is simple if at most two of the $n$ curves intersect in any point: The three diagrams in the top row of Figure 1 are simple, whereas the three diagrams in the bottom row are non-simple. An easy application of Euler’s formula shows that a simple Venn diagram has exactly $2^{n}-2$ crossings, whereas non-simple diagrams may have much fewer. For humans, simple Venn diagrams tend to be easier to read than non-simple diagrams.
@@
An interesting subclass of Venn diagrams are those that can be drawn with $n$ convex curves. This property can be captured purely combinatorially, as follows: A $k$-region in an $n$-Venn diagram is a region that lies inside of exactly $k$ of the $n$ curves (and outside of the remaining nkn-k curves). A monotone Venn diagram is one in which for every $0<k<n$, every $k$-region is adjacent to both a (k1)(k-1)-region and a $(k+1)$-region. Bultena, Grünbaum and Ruskey [BGR99] proved that monotone diagrams are precisely the ones that can be drawn such that every curve has a convex shape. Formally, every diagram made of convex curves is monotone, and every monotone diagram can be convexified. Monotone diagrams can be realized as a wire diagram, where the $n$ curves are drawn as horizontal lines at heights $1,2,\ldots,n$, and each crossing between neighboring curves occurs within a narrow vertical strip. Furthermore, the diagram implicitly wraps around cyclically at the left and right boundary; see Figure 2. Note that a convex realization can be obtained from the wire diagram by placing it in a very thin annulus with large enough radius.
Figure 2. Representation of the 5-Venn diagram from Figure 1 (c) as a wire diagram.
@@ 1.1. Winkler’s conjecture
He reiterated the problem in his ‘Puzzled’ column [WIN12] in the Communications of the ACM. The conjecture is also listed as the first open problem in Ruskey and Weston’s inspiring survey [RW97] on Venn diagrams.
Concerning small cases, the conjecture is easily verifiable for all simple Venn diagrams on $n\in\{3,4,5\}$ curves, simply because there are so few of them. Another class of Venn diagrams for which the conjecture is easily seen to be true are reducible diagrams. An $n$-Venn diagram is reducible if one of its curves can be removed so that the remaining n1n-1 curves form an (n1)(n-1)-Venn diagram. Otherwise the diagram is called irreducible. The diagram in Figure 1 (c) is irreducible, because any 4 of the 5 curves create 20 regions instead of 16. If a diagram is reducible, then the curve $C$ that can be removed splits every region of the resulting (n1)(n-1)-Venn diagram into two, and therefore this curve touches every region in the $n$-Venn diagram. We can thus add an (n+1)(n+1)st curve by following the curve $C$ and subdividing every region into two, changing sides along $C$ exactly once along every segment between two consecutive crossings. Thus, the interesting instances of Conjecture 1 are irreducible diagrams.
@@ 1.1.1. Translation to a Hamilton cycle problem
Figure 3. (a) A (monotone) 4-Venn diagram and its dual graph, a (monotone) 4-Venn quadrangulation; (b) a Hamilton cycle in the quadrangulation; (c) the 5-Venn diagram obtained from (a) by adding the curve corresponding to the Hamilton cycle in (b).
@@
    It is a connected subgraph of $Q_{n}$ that is also spanning, i.e., all $2^{n}$ vertices are present. Specifically, the vertex in $Q(D)$ corresponding to a region of the diagram $D$ is the characteristic vector of the inside/outside relation with respect to each curve, where 0 and 1 represent outside and inside, respectively. [^1] When considering the diagram in the plane, then the inside and outside of a curve $C$ are the bounded and unbounded region of $\mathbb{R}^{2}\setminus C$, respectively.
@@
We refer to an $n$-Venn quadrangulation as monotone, if it satisfies the following additional condition, where the weight $w(x)$ of a bitstring $x$ is the number of 1s in $x$:
  1. 3’
    Every vertex $x$ with weight $0<w(x)<n$ has a neighbor of weight w(x)1w(x)-1 and a neighbor of weight $w(x)+1$.
Monotone $n$-Venn quadrangulations are precisely the duals of monotone $n$-Venn diagrams. Note that 3’ implies 3, as connectedness in a monotone Venn quadrangulation is ensured via the vertices $0^{n}$ and $1^{n}$. Monotone $n$-Venn quadrangulations are special cases of so-called rhombic strips recently studied in [ACF+25, BBM+25].
Extending a given diagram $D$ by a single curve means that the new curve has to enter and leave every region of $D$ exactly once, splitting it into two, one of them inside and the other one outside the newly added curve. This corresponds to traversing a Hamilton cycle in the dual graph $Q(D)$, i.e., a cycle in the graph that visits every vertex exactly once; see Figure 3 (b)+(c).
@@ 1.1.2. The counterexamples
Figure 4. (Left) A counterexample to Winkler’s conjecture for $n=6$, namely a simple (non-monotone) 6-Venn diagram that cannot be extended to a simple 7-Venn diagram by adding a suitable curve. Highlighted are 12 red regions that are surrounded by only 11 blue regions, which is a local obstruction for extendability. [^3] This drawing and the one in Figure 5 were produced using a SageMath program [SCH24] developed by Manfred Scheucher for his and Felsner’s work [FS21] on pseudocircle arrangements, which we used with Scheucher’s permission. (Right) The dual graph of the diagram on the left, namely a 6-Venn quadrangulation $P$ that has no perfect matching and hence no Hamilton cycle. Edges are colored according to the colors of the curves. The highlighted subgraph $H$ contains a subset of 12 red vertices in one partition class and their 11 blue neighbors, witnessing the non-existence of a perfect matching. The green subgraph $L$ is the ladder used for extending the counterexample to all values of $n>6$, and all those extensions contain a copy of $H$.
@@
One may wonder whether Winkler’s conjecture can be salvaged for monotone diagrams, i.e., those that can be drawn with convex curves. Our next theorem shows that this is not possible. A counterexample for $n=7$ is shown in Figure 5.
Theorem 3.
For every $n\geq 7$ there is a monotone $n$-Venn quadrangulation that has no perfect matching and hence no Hamilton cycle. Its dual is a simple monotone $n$-Venn diagram that cannot be extended to a simple $(n+1)$-Venn diagram by adding a suitable curve.
Figure 5. A simple monotone 7-Venn diagram that cannot be extended to a simple 8-Venn diagram by adding a suitable curve. The bottom picture is the same diagram in a wire representation. The subgraph $H$ of the dual Venn quadrangulation is the same as in Figure 3. It corresponds to the highlighted 12 red regions that are surrounded by only 11 blue regions, which is a local obstruction for extendability. The green subgraph $L$ is the ladder used for extending the counterexample to all values of $n>7$.
As mentioned before, for $n\leq 5$ all simple $n$-Venn diagrams are extendable, so Theorem 2 is best possible in the sense that the counterexamples for $n=6$ are minimal. In fact, as part of our proof of Theorem 2 we computed all distinct 6-Venn diagrams and analyzed their properties, summarized in the next theorem, Figure 6 and Table 1. As it turns out, all 72 non-extendable diagrams for $n=6$ are non-monotone, and so Theorem 3 is also best possible.
For the counting results stated in the next theorem, we consider two diagrams the same if they differ only by mirroring and/or stereographic projection. Equivalently, we count their dual $n$-Venn quadrangulations up to graph isomorphism. Prior to this work, the number of simple $n$-Venn diagrams was only known up to $n\leq 5$. Specifically, for $n\leq 4$ there is only one, and for $n=5$ there are 20 [HP97, CHP00]. An exposed diagram is one in which the outer region is bounded by all $n$ curves. It is not hard to see that any monotone diagram is also exposed. Chilakamarri, Hamburger and Pippert [CHP96b] constructed $n$-Venn diagrams for every $n$ in which all regions are (combinatorial) triangles, quadrangles or pentagons, in particular, they are not exposed for $n\geq 6$.
@@
The 6-Venn quadrangulation from Figure 3 is one of the 17 we found that have no perfect matching. The monotone 7-Venn quadrangulation from Figure 5 was also found with computer help, but using SAT solvers instead, prescribing the occurrence of a copy of the ‘obstacle’ subgraph $H$ as one of the constraints. We feel that it is very much out of reach to determine and/or count all simple (monotone) 7-Venn diagrams.
The programs we used for computing all simple 6-Venn diagrams and the result files with all 6-Venn quadrangulations, as well as the monotone 7-Venn quadrangulation corresponding to the diagram in Figure 5, are available for download [14].
@@ 2. Proof of Theorems 2 and 3
Consider an $n$-Venn quadrangulation $P$ that contains a ladder $L$ and a subgraph $H$ with $L\cap H=\emptyset$. Then there is an $(n+1)$-Venn quadrangulation $P^{\prime}$ that contains a ladder $L^{\prime}$ and a copy $H^{\prime}$ of $H$ such that $L^{\prime}\cap H^{\prime}=\emptyset$. If $P$ is monotone, then $P^{\prime}$ is also monotone.
@@
Let $x_{1},\ldots,x_{n}$ and $y_{1},\ldots,y_{n}$ be the vertices of the ladder $L$ in $P$ as defined before, and let $k\in[n]$ be the type of the edges $(x_{i},y_{i})$ for $i=1,\ldots,n$. We write $C$ for the outer cycle of $L$, i.e., for the graph $C:=L\setminus\{(x_{i},y_{i})\mid i=2,\ldots,n-1\}$. Note that in every 4-cycle of a Venn quadrangulation, there are edges of exactly two different types, and opposite edges have the same type. Consequently, both edges $(x_{i},x_{i+1})$ and $(y_{i},y_{i+1})$ of $L$ (and $C$) have the same type for $i=1,\ldots,n-1$, and because of the condition $x_{1}=0^{n}$ and $y_{n}=1^{n}$, every type $j\in[n]\setminus\{k\}$ appears for exactly one pair of opposite edges $(x_{i},x_{i+1})$ and $(y_{i},y_{i+1})$ for some $i=1,\ldots,n-1$. We write $C_{j}^{0}$ and $C_{j}^{1}$ for the two connected subgraphs of $C$ obtained from $C$ by removing both edges $(x_{i},x_{i+1})$ and $(y_{i},y_{i+1})$, such that the first one contains the vertex $x_{1}=0^{n}$ and the second one contains the vertex $y_{n}=1^{n}$.
We assume w.l.o.g. that the 4-cycle $(x_{1},x_{2},y_{2},y_{1})$ bounds the outer face of $P$. Let $P^{-}$ be the graph obtained from $P$ by removing the edges $(x_{i},y_{i})$ of the ladder $L$ for all $i=2,\ldots,n-1$. Let $\operatorname{rev}(P^{-})$ denote the plane graph obtained from $P^{-}$ by mirroring. Appending a 0-bit to all vertices of $P^{-}$ and a 1-bit to all vertices of $\operatorname{rev}(P^{-})$ gives the graphs $P^{-}0$ and $\operatorname{rev}(P^{-})1$. We let $P^{\prime}$ be the plane graph obtained by connecting $P^{-}0$ and $\operatorname{rev}(P^{-})1$ via $2n$ additional type-$(n+1)$ edges $(x_{j}0,x_{j}1)$ and $(y_{j}0,y_{j}1)$ for all $j=1,\ldots,n$. It is straightforward to check that $P^{\prime}$ satisfies conditions 13, and that if $P$ satisfies condition  3’, then $P^{\prime}$ also satisfies  3’. In particular, the connectivity condition  3 is ensured via the subgraphs $C_{j}^{0}$ and $C_{j}^{1}$ for all bit positions $j\in[n]\setminus\{k\}$, and for the bit position $k$ it is ensured via the subgraphs $C0$ and $C1$.
@@
The proof of Theorem 3 works very similarly.
Proof of Theorem 3.
The monotone 7-Venn quadrangulation $P$ dual to the diagram shown in Figure 5 contains a copy of the same subgraph $H$ as considered in the previous proof. As $H$ has no perfect matching, $P$ has no perfect matching either.
Repeatedly applying Lemma 8 to $P$ with the subgraph $H$ and the ladder $L$ marked in Figure 5 yields a monotone $n$-Venn quadrangulation without perfect matching for all $n>7$. This completes the proof. ∎
@@ 3. Proof of Theorem 4
In this section we describe our C++ and SageMath programs that compute all 6-Venn quadrangulations. For the reader’s convenience, the code and data produced by them are available for download [14]. Specifically, the 6-Venn quadrangulations are provided in graph6 format, and in a compact binary adjacency list representation. The following explanations are valid for general $n$, but have a reasonable computation time only for $n\leq 6$. Table 1 provides the counts of Venn diagrams with various properties for all $n=1,\ldots,6$.
@@
Maybe surprisingly, the computational proof of Theorem 4 was the technically most demanding part of this work, even though in the end there are relatively few 6-Venn quadrangulations in total ($10^{7}$ many). In fact, we had first found a counterexample to Winkler’s conjecture for $n=7$ (using different methods that involved SAT solvers) and a proof of Theorem 2 for $n\geq 7$, before finding the few counterexamples for $n=6$.
@@ 3.1. Dynamic programming on two halves of a Venn diagram
    It is a connected subgraph of $Q_{n-1}$ that is also spanning, i.e., all $2^{n-1}$ vertices are present.
@@ 3.1.1. Step (i): Computing all cycles $C$ of $Q_{n-1}$
In the following, we describe another observation that allows us to restrict the lengths of the boundary cycles to consider. As mentioned before, the total number of crossings in a simple $n$-Venn diagram $D$ is $2^{n}-2$. Clearly, this number equals the number of faces of $Q(D)$. The number of vertices of $Q(D)$ is $2^{n}$, and therefore Euler’s formula yields $2^{n+1}-4$ for the number of edges of $Q(D)$. These edges split into $n$ different matchings $M_{1},M_{2},\ldots,M_{n}$ according to their types, and we can therefore always guarantee that one matching $M_{i}$ has size $|M_{i}|\geq\frac{2^{n+1}-4}{n}$, and the fraction on the right should be rounded up to the next even integer, because the corresponding boundary cycle in the hypercube must have even integral length. For $n=6$ this yields the lower bound $|M_{i}|\geq 22$ for one of the matchings. By symmetry, we can thus restrict our search to cycles $C$ in $Q_{5}$ of lengths $22,24,26,28,30,32$. Note that if $C$ has the maximum length $2^{n-1}$, then it is a Hamilton cycle in $Q_{n-1}$, and the resulting $n$-Venn diagram $H$ obtained by pairing two compatible $C$-quadrangulations is reducible and thus extendable.